# Batayan: A Filipino NLP benchmark for evaluating Large Language Models

### v1

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities on widely benchmarked high-resource languages; however, linguistic nuances of under-resourced languages remain unexplored. We introduce Batayan, a holistic Filipino benchmark designed to systematically evaluate LLMs across three key natural language processing (NLP) competencies: understanding, reasoning, and generation. Batayan consolidates eight tasks, covering both Tagalog and code-switched Taglish utterances. Our rigorous, native-speaker-driven annotation process ensures fluency and authenticity to the complex morphological and syntactic structures of Filipino, alleviating a pervasive translationese bias in existing Filipino corpora. We report empirical results on a variety of multilingual LLMs, highlighting significant performance gaps that signal the under-representation of Filipino in pretraining corpora, the unique hurdles in modeling Filipino’s rich morphology and construction, and the importance of explicit Filipino language support and instruction tuning. Moreover, we discuss the practical challenges encountered in dataset construction and propose principled solutions for building culturally and linguistically-faithful resources in under-represented languages. We also provide a public benchmark and leaderboard as a clear foundation for iterative, community-driven progress in Filipino NLP.
@@ 1 Introduction
Filipino, despite being the national language of the Philippines and being spoken by over 80 million people, remains an under-resourced language. It is under-represented in multilingual LLMs, and existing corpora are domain-specific, non-multilingual Cajote et al. (2024), and created by non-native speakers Quakenbush (2005); Dita et al. (2009). Specifically for LLM benchmarks, Filipino is either excluded from multilingual benchmarks, suffer from limited task diversity, or have serious deficiencies in grammatical correctness, completeness, and diversity Bandarkar et al. (2024); NLLB Team et al. (2022).
@@
Our threefold contributions attempt to address these challenges. First, we present Batayan, [^1] The word “batayan” means “basis” or “groundtruth”. a holistic Filipino benchmark for evaluating large language models across 8 distinct tasks spanning natural language understanding, reasoning, and generation. We place a rigorous emphasis on authenticity to natural Filipino language use through native speaker translation and annotation, addressing the limitations of existing Filipino datasets. Batayan is released as part of SEA-HELM [^2] https://leaderboard.sea-lion.ai/, a leaderboard for comprehensive evaluation of LLMs across linguistic and reasoning tasks for Southeast Asian languages. Second, we provide extensive evaluation results from testing 8 prominent LLMs on Batayan, revealing significant disparities in model performance across different linguistic capabilities in Filipino. Third, we document systematic challenges and methodological considerations in creating high-quality Filipino NLP datasets, particularly highlighting issues in translation fluency, vocabulary adaptation, and the preservation of Filipino’s rich morphological features. These insights provide a practical framework for future development of Filipino language resources.
@@ 2.1 Language of Evaluation
While Filipino is the official language of the Philippines, it has been argued in literature that its de facto lingua franca is Taglish, the practice of code-switching between English and Tagalog Go and Gustilo (2013). Naturally-occurring datasets in Filipino (e.g., those mined from social media, recorded interviews) contain some code-switching Bautista (2004). We defer the problematization of the debated difference between Tagalog and Filipino, and use the two terms interchangeably.
Standard Tagalog, while text-rich, is considered under-resourced, lacking the linguistic data, tools, and resources for effective natural language processing Miranda (2023b). To address data scarcity, developers take advantage of high-resource languages such as English through translations Goyal et al. (2022); Doddapaneni et al. (2023). Furthermore, leveraging multilingual datasets is potent not only because it enhances the performance of models trained on limited resources but also because code-switching and bilingualism is a common phenomena in the Philippines Tupas and Martin (2017).
Taglish is characterized to possess communicative efficiency. Compared to Standard Tagalog or English, Taglish, provides a more convenient way of communicating a message Bautista (2004). For this consideration, application of NLP becomes more helpful as users are more likely to interact freely when using a language familiar to them.
@@ 2.2 Language Design Principles
The design of Batayan adhered to principles that emphasize linguistic authenticity and representativeness in word choice, sentence structure, and grammar.
In terms of language use and word choice, we employed common Filipino and Taglish vocabulary, balancing loanwords with native Filipino terms by prioritizing colloquial usage. This was also done when words had both English and Filipino variants, taking into account spelling and orthographic preferences prevalent among native speakers.
Regarding sentence structure, we prioritized sentences with attention to natural syntax and the choice of ayos (sentence arrangement), namely direct (karaniwang ayos, KA, lit. usual order) or inverted (di-karaniwang ayos, DKA, lit. unusual order) forms Tanawan et al. (2008). In KA, Filipino constructions follow the typical predicate-initial word order Malicsi (2013). In contrast, DKA, which is also referred to linguistically as an ay-inversion, is a type of construction in Filipino where non-predicative constituents (such as simuno, lit. grammatical subject) are “fronted” or shifted to precede the predicate, “marking” them as the “topic” of the sentence Kroeger (1993). Inverted constructions are often used in formal settings, and could therefore be deemed unnatural in most situations Pizarro-Guevara (2010). This characteristic of Filipino is of interest due to our observation of automatic translations preferring DKA (see Section 4.1); hence, the further need for native re-translations with a preference for KA. The distinction between KA and DKA with respect to natural syntax and translation style is further discussed in the same section.
For translated datasets, compliance with Filipino grammar was also a critical aspect of the dataset design, with adherence to grammatical guidelines set by the Komisyon sa Wikang Filipino Almario (2014). In translation, we followed principles aimed at preserving the source material’s intent while adjusting for cultural and contextual relevance as suggested by Almario (2016). Translations were performed by native Filipino speakers in the identified lingua franca, whenever appropriate, to maintain semantic equivalence without resorting to overly literal phrasing.
@@ 3.1 Task Selection
Previous work on Filipino language model evaluation has been largely fragmented. While prior studies have made important contributions in assessing large-scale language models Cruz and Cheng (2020, 2022), and various researchers have developed task-specific datasets for named entity recognition Miranda (2023a), sentiment analysis Villavicencio et al. (2021), and other isolated tasks, the field lacks a unified, comprehensive benchmark for systematic evaluation of model capabilities across different linguistic dimensions.
Through BATAYAN, we aim to significantly expand the scope of Filipino language evaluation by introducing tasks that assess a gamut of linguistic capabilities. Specifically, we focus on integrating more challenging Natural Language Understanding (NLU) tasks alongside Natural Language Reasoning (NLR) tasks, while introducing novel Natural Language Generation (NLG) tasks previously unexplored in the Filipino context. This comprehensive approach allows for a more thorough assessment of models’ ability in Filipino given its rich morphology and unique linguistic characteristics.
Our benchmark design and selection of tasks are informed by established multilingual evaluation frameworks, particularly BHASA Leong et al. (2023), which provides systematic evaluation across a number of Southeast Asian languages. We also draw insights from widely-adopted language benchmarks such as the XTREME multilingual benchmark Hu et al. (2020) and the IndoNLU Indonesian benchmark Wilie et al. (2020), which have demonstrated the importance of testing various aspects of linguistic competence.
For this study, we carefully selected eight tasks that span three key competencies of NLP. For NLU: Paraphrase Identification (PI), Question Answering (QA), Sentiment Analysis (SA), Toxicity Detection (TD). For NLR: Causal Reasoning (CR), Natural Language Inference (NLI). For NLG: Abstractive Summarization (AS), Machine Translation (MT). These tasks are detailed in Appendix B.
3.2 Dataset Collection and Annotation
Competency Task Dataset Source Language Output (Options) Domain Adaptation Applied
NLU PI PAWS Zhang et al. (2019) English label (2) Wikipedia Native translation
QA Belebele Bandarkar et al. (2024) Filipino span (4) Wikinews Native translation
SA PH Elections Cabasag et al. (2019) Taglish label (3) Social media None
TD PH Elections Cabasag et al. (2019) Taglish label (2) Social media None
NLR CR Balanced COPA Kavumba et al. (2019) English label (2) General Native translation
NLI XNLI Conneau et al. (2018) English label (3) General Native translation
NLG AS XL-Sum Hasan et al. (2021) English summary News Native translation
MT FLORES 200 NLLB Team et al. (2022) Filipino translation Wikinews Native translation
Table 1: Source datasets for each task in Batayan.
@@ 3.2 Dataset Collection and Annotation
To construct Batayan, we curated open-source corpora with clear provenance when possible, prioritizing datasets that exhibit authentic language use across various domains such as social media, news articles, and other publicly available texts. We identified existing Filipino-language datasets for MT, QA, and TD. For the SA task, we chose to repurpose an existing dataset and annotate them for the aforementioned task.
Tasks such as AS, CR, NLI, and PI lacked a native Filipino corpus. As such, we first identified existing English corpora, generated initial translations using automatatic translation tools, then conducted a manual review and revision to generate high-quality and fluent translations. The authors [^3] Demographics: working professionals and university students, aged between 21–30 years old, native Filipino speakers, lived in the Philippines for a majority of their lives. acted as annotators and raters for this study. They were grouped into teams of three and performed translations and reviews, focusing on cultural and linguistic relevance to ensure that the translated texts resonate with Filipino contexts.
For each task, we randomly sampled $5n$ entries for a target size of $n$, maintaining balanced class distributions. Entries were filtered by length (20–2000 characters) and quality ($\leq$50grammatical errors). Afterwards, each sample in each task underwent evaluation by teams of three against three criteria: completeness, fluency, and sensibility. Samples were assessed as complete only if they contained well-formed sentences, excluding standalone dependent clauses or fragmentary headers and titles lacking complete meaning. Fluency was evaluated on native-like constructions based on multiple factors including appropriate verb conjugations (e.g., avoiding constructions like “*Nagpapasalamatan ako sa iyo”, lit. “*I give thanks to you.”), natural word choices (avoiding awkward phrases like “*Si Pacquiao ang kamao ng bansa” instead of “Si Pacquiao ang pambansang kamao”, in English: “Pacquiao is the nation’s fist”), and preference for KA sentence structure where contextually appropriate. Samples were assessed as sensible if relevant to the task, screening for confounding factors that would not effectively test the intended language capability (e.g., incorrect answers in QA mislabeled as correct).
For AS, we employed additional criteria based on previous work Leong et al. (2023) to assess the appropriateness of the summary in reference to the content of the provided passage. We score samples on faithfulness (binary), relevance (0–3), fluency (0–3), and coherence (0–3), with three independent raters per entry. For faithfulness, raters evaluated the factual consistency between the summary and source article, with particular attention to fact preservation and penalization of hallucinated content. Only summaries receiving unanimous 1/1 faithfulness scores from all raters were considered for inclusion, ensuring strict maintenance of factual integrity. For relevance, raters evaluated both coverage of essential information and information filtering, penalizing summaries containing redundancies or excess details. For fluency, raters assessed for proper formatting, appropriate capitalization, and grammatical construction, with higher scores awarded to summaries exhibiting natural Filipino language patterns. For coherence, raters assessed how effectively information flowed between sentences, with higher scores given to summaries that built logically from sentence to sentence to create a cohesive narrative about the topic.
Only samples passing all applicable criteria were included in the final dataset. In cases where initial sampling yielded insufficient qualifying entries, deficient samples were carefully corrected by authors through targeted improvements to grammar, translation of English passages, or other identified issues while maintaining authentic Filipino language patterns.
Joint agreement on our evaluation criteria is presented in Appendix A. This process prioritized agreement on fundamental classification criteria rather than traditional metrics which use chance correction like Cohen’s kappa which are less suitable for binary categorical tasks Powers (2012) and classification tasks in general Delgado and Tibau (2019); Feng and Zhao (2016).
3.3 Source Datasets
@@
Table 1 outlines key attributes of each dataset, such as language, output type, domain, and any adaptations applied to ensure alignment with natural Filipino language construction. More detailed statistics can be found in Appendix B.
Native review and translation was applied for all tasks (except for SA and TD) to enhance contextual and linguistic accuracy. We chose to maintain the naturally-occurring language for sentiment analysis and toxicity detection to capture the nuances of code-switched Taglish usage prevalent in social media in the Philippines.
For the SA task, three native Filipino speakers reannotated the Philippine election-related tweets Cabasag et al. (2019) dataset with three sentiment polarities: positive, negative, and neutral. Inter-annotator agreement was calculated, which yielded a Cohen’s kappa of 0.8202 and a Krippendorf’s alpha of 0.8268, indicating substantial agreement.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities on widely benchmarked high-resource languages. However, linguistic nuances of under-resourced languages remain unexplored. We introduce Batayan, a holistic Filipino benchmark that systematically evaluates LLMs across three key natural language processing (NLP) competencies: understanding, reasoning, and generation. Batayan consolidates eight tasks, three of which have not existed prior for Filipino corpora, covering both Tagalog and code-switched Taglish utterances. Our rigorous, native-speaker-driven adaptation and validation processes ensures fluency and authenticity to the complex morphological and syntactic structures of Filipino, alleviating the pervasive translationese bias in existing Filipino corpora. We report empirical results on a variety of open-source and commercial LLMs, highlighting significant performance gaps that signal the under-representation of Filipino in pre-training corpora, the unique hurdles in modeling Filipino’s rich morphology and construction, and the importance of explicit Filipino language support. Moreover, we discuss the practical challenges encountered in dataset construction and propose principled solutions for building culturally and linguistically-faithful resources in under-represented languages. We also provide a public evaluation suite as a clear foundation for iterative, community-driven progress in Filipino NLP.
@@ 1 Introduction
Filipino, despite being the national language of the Philippines and being spoken by over 80 million people, remains an under-resourced language. It is under-represented in multilingual LLMs, and existing corpora are domain-specific, non-multilingual Cajote et al. (2024), and created by non-native speakers Quakenbush (2005); Dita et al. (2009). Filipino is often excluded from multilingual LLM benchmarks, suffer from limited task diversity, or have serious deficiencies in grammatical correctness, completeness, and diversity Bandarkar et al. (2024).
@@
Our threefold contributions attempt to address these challenges. First, we present Batayan, [^1] The word “batayan” means “basis” or “ground truth”. a holistic Filipino benchmark for evaluating LLMs across 8 distinct natural language processing (NLP) tasks spanning understanding, reasoning, and generation. We place a rigorous emphasis on authenticity to natural Filipino language use through native speaker translation and annotation, addressing translationese bias and other limitations of existing Filipino datasets. Batayan is released as part of SEA-HELM, a publicly available evaluation suite [^2] https://github.com/aisingapore/sea-helm and leaderboard [^3] https://leaderboard.sea-lion.ai for assessing LLMs across linguistic and reasoning tasks for Southeast Asian languages. Second, we provide extensive evaluation results from a comprehensive set of open-source and commercial LLMs on Batayan, from differing sizes (7B–671B parameters) and Filipino-language representation, revealing significant disparities in model performance across different linguistic capabilities in Filipino. Third, we document systematic challenges and methodological considerations in creating high-quality Filipino NLP datasets, particularly highlighting issues in translation fluency, vocabulary adaptation, and the preservation of Filipino’s rich morphological features. These insights provide a practical framework for future development of Filipino language resources.
@@ 2 Considerations for Filipino Evaluations
Comp. Task Dataset # test samples Target1 Language Source2 Adaptation Our contribution
NLU PI PAWS Zhang et al. (2019) 2,000 label (2) English English-sourced Native translation
QA Belebele Bandarkar et al. (2024) 900 span (4) Filipino English-adapted Native re-translation
SA PH Elections Cabasag et al. (2019) 5,160 label (3) Taglish Natively-sourced No adaptation
TD PH Elections Cabasag et al. (2019) 5,160 label (2) Taglish Natively-sourced No adaptation
NLR CR Balanced COPA Kavumba et al. (2019) 500 label (2) English English-sourced Native translation
NLI XNLI Conneau et al. (2018) 5,010 label (3) English English-sourced Native translation
NLG AS XL-Sum Hasan et al. (2021) 11,535 summary English English-sourced Native translation
MT FLORES 200 NLLB Team et al. (2022) 1,012 translation Filipino English-adapted Native re-translation
  • 1
    Number of options are shown in parenthesis
  • 2
    English-adapted: previously translated from English; English-sourced: originally in English; Natively-sourced: originally in Taglish
Table 1: Source datasets for each task in Batayan.
@@ 2.1 Language of Evaluation
While Filipino is the official language of the Philippines, it has been argued in literature that the de facto lingua franca is Taglish, the practice of code-switching between English and Tagalog Go and Gustilo (2013). Naturally-occurring datasets in Filipino (e.g., mined from social media, recorded interviews) typically contain some code-switching Bautista (2004). We defer the problematization of the debated difference between Tagalog and Filipino, and use the two terms interchangeably.
Filipino, while text-rich, is considered under-resourced, lacking the linguistic data, tools, and resources for effective natural language processing Miranda (2023b). To address data scarcity, developers take advantage of high-resource languages such as English through translation Goyal et al. (2022); Doddapaneni et al. (2023). Furthermore, leveraging multilingual datasets is potent not only because it enhances the performance of models trained on limited resources, but also because code-switching and bilingualism is a common phenomenon in the Philippines Tupas and Martin (2017).
@@ 2.2 Language Design Principles
The design of Batayan emphasized linguistic authenticity and representativeness in word choice, sentence structure, and grammar. In terms of language use and word choice, we employed common Filipino and Taglish vocabulary, balancing loanwords with native Filipino terms by prioritizing colloquial usage. This was also done when words had both English and Filipino variants, taking into account spelling and orthographic preferences prevalent among native speakers.
Regarding sentence structure, we prioritized sentences with attention to natural syntax and the choice of ayos (sentence arrangement), namely direct (karaniwang ayos, KA, lit. usual order) or inverted (di-karaniwang ayos, DKA, lit. unusual order) forms Tanawan et al. (2008). In KA, Filipino constructions follow the typical predicate-initial word order Malicsi (2013). In contrast, DKA, which is also referred to linguistically as an ay-inversion, is a type of construction in Filipino where non-predicative constituents (such as simuno, lit. grammatical subject) are “fronted” or shifted to precede the predicate, marking them as the “topic” of the sentence Kroeger (1993). Inverted constructions are often used in formal settings, and could therefore be deemed unnatural in most situations Pizarro-Guevara (2010). This characteristic of Filipino is of interest due to our observation of machine translations preferring DKA (see Section 4.1); hence, the further need for native re-translations with a preference for KA.
2.3 Related Datasets and Benchmarks
Previous work on Filipino language model evaluation has been largely fragmented. Researchers have made efforts in developing datasets for Filipino language tasks such as fake news detection Cruz et al. (2020), named entity recognition Miranda (2023a), natural language inference Cruz et al. (2021), sentiment analysis Villavicencio et al. (2021), and others. However, the field lacks a unified, comprehensive benchmark for systematic evaluation of LLMs across different linguistic dimensions in Filipino.
Unlike existing Southeast Asian benchmarks that either overlook Filipino or treat it through machine translation, Batayan is the first holistic, native-speaker-driven evaluation suite tailored specifically for Filipino NLP. For instance, BHASA provides a multilingual SEA benchmark covering Indonesian, Tamil, Thai, and Vietnamese, but it does not include Filipino Leong et al. (2023). Similarly, IndoNLU focuses exclusively on Indonesian understanding tasks and lacks any generative or reasoning evaluations Wilie et al. (2020). SeaExam and SeaBench introduce local exam and open-ended Vietnamese, Thai, and Indonesian queries but exclude Filipino entirely Liu et al. (2025), and SailCompass evaluates robustness for SEA languages without any Filipino components Guo et al. (2024). SeaEval provides a broad multilingual and multicultural benchmark across 29 datasets for Southeast Asian languages, including Filipino, but still relies primarily on machine translations and generic multilingual prompts Wang et al. (2024).
In contrast, Batayan integrates eight tasks—including three novel Filipino-centric ones—grounded in native translations and rigorous acceptability judgments, ensuring fluency, authenticity, and sensitivity to Filipino’s complex morphological and code-switching patterns.
@@ 3.1 Task Selection
Through BATAYAN, we significantly expand the scope of Filipino LLM evaluation by introducing tasks that assess a gamut of NLP competencies: understanding (NLU), generation (NLG), and reasoning (NLR). Our Filipino benchmark incorporates machine translation (MT), natural language inference (NLI), question answering (QA), sentiment analysis (SA), and toxicity detection (TD). We also release three novel tasks that have not been previously explored in the Filipino context, namely abstractive summarization (AS), causal reasoning (CR), and paraphrase identification (PI).
Our benchmark design and selection of tasks are informed by established multilingual evaluation frameworks, particularly BHASA Leong et al. (2023) and XTREME Hu et al. (2020).
3.2 Dataset Selection
@@
To construct Batayan, we curated open-source corpora with clear provenance when possible, prioritizing datasets that exhibit authentic language use across various domains such as social media, news articles, and other publicly available texts. We identified existing Filipino-language datasets for MT, QA, SA, and TD. Meanwhile, the AS, CR, NLI, and PI tasks lacked an expert-annotated, natively-sourced Filipino corpus. Given the scarcity of high-quality Filipino corpora for these tasks, we chose to natively translate and adapt existing English corpora into Filipino following the methods detailed in Section 3.3. In translation, we deliberately “naturalized” the language use by preferring KA, colloquial terms, and other adaptations to preserve native Filipino speech.
@@
Table 1 outlines key attributes of each dataset, such as output type, domain, language, source, and any adaptations that we applied to ensure alignment with natural Filipino language construction. More detailed statistics can be found in Appendix A.
3.3 Dataset Adaptation and Annotation
Figure 2: Dataset curation process for Batayan.
For each task, entries were filtered by length (20–2000 characters), then randomly sampled $5n$ entries for a target size of $n$, maintaining balanced class distributions.
Native speaker acceptability judgments for quality filtering. For tasks that were natively-sourced (such as TD) or were previously translated from English to Filipino (such as MT, QA), we filtered entries by text quality. Our evaluation of language quality was driven by acceptability judgments from native Filipino speakers [^4] Demographics: Working professionals, research faculty, and university students, aged between 21–30 years old. Native bilingual speakers and code-switchers of Filipino and English. Lived in the Philippines for a majority of their lives. in groups of three so that context and natural language variation are appropriately considered. Relying on standardized automated metrics would be less effective for languages with rich morphological and syntactic traits Rivera-Trigueros (2022) such as Filipino.
Annotation for SA task. We chose to repurpose the natively-sourced TD dataset due to the authenticity and diversity of registers in the dataset. Three native speakers annotated the Philippine election-related tweets Cabasag et al. (2019) dataset for the SA task with three sentiment polarities: positive, negative, and neutral. Inter-annotator agreement was calculated using Cohen’s kappa Cohen (1960), which is the difference between observed agreement and the probability of chance agreement over the probability that the raters did not agree by chance, and Krippendorff’s alpha Krippendorff (2018), which is 1 minus the ratio of observed disagreement by raters and expected chance disagreement). The resulting Cohen’s kappa of 0.8202 and Krippendorf’s alpha of 0.8268 indicate substantial agreement. For both TD and SA tasks, we chose to maintain the naturally-occurring language to capture the nuances of code-switched Taglish usage prevalent in social media in the Philippines.
Authoritative translation guidelines. For datasets that were adapted into Filipino, a critical aspect of dataset design was ensuring compliance with official guidelines set by the Komisyon sa Wikang Filipino (Commission on the Filipino Language) of the Philippines. In translation, we followed prescribed principles aimed at preserving the source material’s intent while adjusting for cultural and contextual relevance Almario (2016) and effective writing Almario (2014). Translations were performed by native Filipino speakers in the identified lingua franca, whenever appropriate, to maintain semantic equivalence without resorting to overly literal phrasing.
Native re-translation of previously translated datasets. We observed that the QA and MT datasets were of subpar quality and suffered from issues with inaccuracies and translationese Gellerstam (1986); Riley et al. (2020) such as the use of unnaturally formal language or inappropriate word choices. As such, these were re-translated to enhance contextual and linguistic accuracy.
Native translation of English datasets. For multilingual tasks that had no equivalent Filipino subsets (such as AS, CR, NLI, and PI), we generated initial translations using Helsinki NLP’s OPUS model. This model was selected over other off-the-shelf tools, such as Google Translate, as it demonstrated superior performance in terms of Filipino translation quality and accuracy as tested by the authors. These translations were then manually revised by native Filipino speakers to generate high-quality and fluent translations. Our native-speaker-driven evaluation process focused on cultural and linguistic relevance to ensure that the translated texts resonate with Filipino contexts.
3.4 Dataset Evaluation
Human evaluation for data quality. Each sample in each dataset underwent further quality assessment by teams of three raters against three binary criteria: completeness, fluency, and sensibility. Samples were assessed as complete only if they contained well-formed sentences, excluding standalone dependent clauses or fragmentary headers and titles lacking complete meaning. Fluency was evaluated on native-like constructions based on multiple factors including appropriate verb conjugations (e.g., avoiding constructions like “*Nagpapasalamatan ako sa iyo”, lit. “*I give thanks to you.”), natural word choices (avoiding awkward phrases like “*Si Pacquiao ang kamao ng bansa” instead of “Si Pacquiao ang pambansang kamao”, lit.: “Pacquiao is the nation’s fist”), and preference for KA sentence structure where contextually appropriate. Samples were assessed as sensible if relevant to the task, screening for confounding factors language capability (e.g., incorrect answers in QA mislabeled as correct or requiring re-translation).
Additional criteria for abstractive summarization. We employed additional criteria based on previous work Leong et al. (2023) to assess the appropriateness of the summary in reference to the content of the provided passage. We scored samples on binary criteria (completeness, fluency, and faithfulness) and a Likert-scaled (0-3) criteria (relevance, fluency, and coherence of the summary), with three independent raters per entry. For faithfulness, raters evaluated the factual consistency between the summary and source article, with particular attention to fact preservation and penalization of hallucinated content. For relevance, raters evaluated both coverage of essential information and information filtering, penalizing summaries containing redundancies or excess details. For fluency, raters assessed for proper formatting, appropriate capitalization, and grammatical construction, with higher scores awarded to summaries exhibiting natural Filipino language patterns. For coherence, raters assessed how effectively information flowed between sentences, with higher scores given to summaries that built logically from sentence to sentence to create a cohesive narrative about the topic.
3.5 Dataset Statistics
Competency Task Label # test samples
NLU PI True 200
False 200
QA span 100
SA Negative (Negatibo) 200
Neutral (Neutral) 200
Positive (Positibo) 200
TD Clean (Malinis) 200
Toxic (Mapoot) 200
NLR CR Cause (Sanhi) 200
Effect (Bunga) 200
NLI Contradiction 200
Entailment 200
Neutral 200
NLG AS summary 100
MT engtgl translation 600
tgleng translation 600
Total 3,800
Table 2: Class distribution per task in Batayan.
For all datasets except AS, only entries that were unanimously rated as demonstrating all 3/3 binary criteria were sampled for the final test splits. For AS, only entries that received a unanimous 3/3 binary criteria assessment and at least and average score of 2 for the scaled criteria were considered for inclusion, ensuring strict maintenance of factual integrity. In cases where initial sampling yielded insufficient qualifying entries, deficient samples were carefully corrected by authors through targeted improvements to grammar, translation of English passages, or other identified issues while maintaining authentic Filipino language patterns.
Joint agreement on our evaluation criteria is presented in Appendix C. We present joint agreement over traditional metrics such as Cohen’s kappa and Krippendorff’s alpha because metrics that use chance correction are less suitable for binary categorical tasks Powers (2012) and classification tasks in general Delgado and Tibau (2019); Feng and Zhao (2016).
Overall, Batayan provides 8 distinct tasks with 3,800 test instances. Table 2 provides the distribution of classes for each dataset, and more detailed statistics can be found in Appendix A. We also provided example splits composed of 5 entries that serve as exemplars for few-shot prompting.